OpenStudy (happykiddo):

How is this factored::::: 125x^3-27=0 using the sum or difference of cubes to find the real or imaginary solutions

OpenStudy (mertsj):

difference of two cubes.

OpenStudy (mertsj):

(5x)^3-(3)^3

OpenStudy (happykiddo):

but what about the part where you plug it into the formula my real question is does the -3 stay negativ when plug-in her into the formula where the -b is

OpenStudy (campbell_st):

the general form is \[a^3 - b^3 = (a -b)(a^2 + ab + b^2)\] in your question a = 5x and b = 3

OpenStudy (happykiddo):

thanks :) i wrote the formula wrong

OpenStudy (campbell_st):

glad to help